OpenStudy (anonymous):

ari’s chickens eat their pellets at a constant rate throughout each day. What is the slope of the line that represents the ounces of pellets eaten? Ari’s Chicken Pellets Days 2 4 6 8 Ounces of Pellets remaining 39 33 27 21

OpenStudy (anonymous):

The general equation for a linear graph is y = ax + b a is the slope a is just the difference between each number, which is -6 as the numbers are decreasing by 6 each time

OpenStudy (texaschic101):

use 2 of your sets of points.... (2,39).....x1 = 2 and y1 = 39 (4,33)...x2 = 4 and y2 = 33 now use the slope formula : (y2 - y1) / (x2 - x1) now sub and solve

OpenStudy (anonymous):

so the slope is as the day gap is 2 so slope = -6/2 = ?